You have $100 dollars. You must purchase a total of 100 animals. Dogs costs $10 each. Cats costs $3 each. Mouse cost .50 cents each. how many Dogs, Cats, Mice do I need to total up to 100 animals and I have to spend the entire $100 dollars.
Answer by ankor@dixie-net.com(22740)

You have $100 dollars. You must purchase a total of 100 animals. Dogs costs $10 each. Cats costs $3 each. Mouse cost .50 cents each. how many Dogs, Cats, Mice do I need to total up to 100 animals and I have to spend the entire $100 dollars.
:
Here's one solution
:
Let d = no. of dogs
Let c = no. of cats
Let m = no. of mice
:
No. of animal equation:
d + c + m = 100
:
Total cost of animals equation
10d + 3c + .5m = 100
:
There can't be very many dogs, 5 at the most
Assume there are 5 dogs
10(5) + 3c + .5m = 100
3c + .5m = 100 - 50
3c + .5m = 50
:
The number of animal equations
5 + c + m = 100
c + m = 95
:
Multiply the $$ equation by 2 and subtract the above equation
6c + 1m = 100
1c + 1m = 95
---------------
5c + 0m = 5
c = 5/5
c = 1 cat
:
Find mice using c + m = 95
1 + m = 95
m = 94 mice
:
Let's assume we have 5 dogs, 1 cat, and 94 mice a total of 100 animals
:
Check the total cost
5(10) + 3(1) + .5(94) =
50 + 3 + 47 = $100
: